# Hydrodynamic flows induced by localized torques (rotlets) in wedge-shaped geometries

### v1

@@ Abstract
Wedge-shaped geometries in low-Reynolds-number flows are of increasing importance, for instance, in the design of microfluidic devices. The corresponding Green’s functions describing the induced flow in response to a locally applied force were derived some time ago. To achieve a complete characterization of particle motion at low Reynolds numbers, we derive the flow response to locally applied torques. This is accomplished through a direct calculation based on the Fourier–Kontorovich–Lebedev transform. We then illustrate the resulting flow fields, highlighting their structure, key features, and dependence on the geometry and orientation of the applied torque. Based on these solutions, we compute the corresponding hydrodynamic mobility tensor that couples torque and motion. Owing to the broken spatial symmetry imposed by the wedge-shaped confinement, a particle subjected to a torque will experience not only rotational motion but also translational motion. These results provide analytical tools relevant for predicting and controlling particle behavior in confined microfluidic environments.
@@ I Introduction
Wedge-shaped geometries have been studied in hydrodynamics because of their peculiar effect on intrinsic fluid flows, and because of their practical abundance. If a net velocity is prescribed on one surface of the wedge in radial direction under low Reynolds number conditions, vortices will be induced near the edge of the wedge, specifically at sufficiently low opening angles of the wedge [1]. For instance, such net motion along one surface can be introduced by radial motion of the surface under no-slip conditions. Wedge-shaped geometries were further investigated under low-Reynolds-number conditions concerning their induced flows due to chemically active surfaces. It was demonstrated that resulting chemical concentration gradients can induce corresponding vortices near the edge of the wedge [2].
@@
Generally, the response to applied higher-order localized force distributions, such as force dipoles or quadrupoles, can be derived from this Green’s function, which is particularly relevant for modeling microswimmers and other active particles [17, 18, 19, 20, 21, 22]. Application of a torque at one point within the fluid is another mode of action on the system. Here, we derive the resulting fluid flow of such a “rotlet” applied to the fluid explicitly. In practice, example situations of imposed rotations at small length scales comprise induced mixing by rotational components [23, 24] or micro-rheological investigations by rotating local probes [25, 26, 27, 28, 29, 30]. Typically, such rotations are induced using magnetic probe particles in reorienting external magnetic fields [31, 32]. In our context of the Green’s function, we would consider them as point-like.
@@ II.1 Governing equations
A general solution to Eq. (1) is given by the Papkovich–Neuber–Imai representation as [34, p. 313],
@@ II.2.1 General overview and definitions
To determine the velocity field generated by a point torque in a wedge-shaped geometry, we employ the Fourier–Kontorovich–Lebedev (FKL) transform. In this approach, the axial $z$, and radial $r$ coordinates, are transformed into their corresponding wavenumbers, denoted by $k$ and $\nu$, respectively. The transform was first introduced by the Soviet mathematicians Kontorovich and Lebedev to address some boundary value problems [35, 36, 37]. The mathematical framework was later developed by Lebedev [38, 39]. We refer the reader to Erdélyi et al. [40, p. 75] for additional details on this integral transform.
@@ III.1 Free-space solution
$f$ $\mathscr{T}_{i\nu}\{f\}$
@@
We next employ the Table of Integral Transforms by Erdélyi et al. [42, p. 175] to derive the FKL transform of $1/s$ as
@@ III.2 Complementary solution
$A_{j}=\frac{2\pi q_{\parallel}}{\nu\operatorname{sh}(\pi\nu)}\left(\frac{\Lambda_{j}}{\rho}+H_{j}\,\frac{\partial}{\partial\rho}\right)K_{i\nu}(|k|\rho)\,,\qquad A_{j}^{\dagger}=\frac{2\pi q_{\parallel}}{\nu\operatorname{sh}(\pi\nu)}\left(\frac{\Lambda_{j}^{\dagger}}{\rho}+H_{j}^{\dagger}\,\frac{\partial}{\partial\rho}\right)K_{i\nu}(|k|\rho)\,$, (24)
which is written in terms of $\Lambda_{j}$, $H_{j}$, $\Lambda_{j}^{\dagger}$, and $H_{j}^{\dagger}$. For transverse torque, and for $j\in\{x,y,w\}$, the coefficients are expressed as
@@
which is written in terms of $\Delta_{j}$ and $\Delta_{j}^{\dagger}$. The reason we include $\nu\operatorname{sh}(\pi)$ in the denominator of these expressions is to simplify the resulting formulas upon performing the inverse transformation; see Eq. (13). In addition, $q_{\parallel}$ and $q_{\perp}$ denote the axial and transverse components of the point torque given in Eq. (4).
@@ III.2.1 Rotlet oriented in the axial direction
In addition, $H_{w}=H_{w}^{\dagger}=0$. The expressions for the remaining coefficients are considerably more complex, so we simplify them using abbreviations. We define the shorthand notations
@@
$\Lambda_{j}$ $=\Gamma_{\sigma}\,\nu\left(a_{j}\nu\sin(2\alpha)+b_{j}\right)\,,\qquad\quad H_{j}=\Gamma_{\sigma}\left(p_{j}\nu\sin(2\alpha)+w_{j}\right)\,$, (32a)
$\Lambda_{j}^{\dagger}$ $=\Gamma_{-\sigma}\,\nu\left(a_{j}^{\dagger}\nu\sin(2\alpha)+b_{j}^{\dagger}\right)\,,\quad\quad H_{j}^{\dagger}=\Gamma_{-\sigma}\left(p_{j}^{\dagger}\nu\sin(2\alpha)+w_{j}^{\dagger}\right)\,$, (32b)
where $\sigma=-1$ for $j=x$ and $\sigma=1$ for $j=y$. The coefficients defining $\Lambda_{j}$ and $\Lambda_{j}^{\dagger}$ are obtained as
@@
The parameters defining $H_{j}$ and $H_{j}^{\dagger}$ are obtained as
@@
Figure 2: Streamlines and quiver plots of the viscous flow field induced by an axially oriented rotlet singularity in a wedge-like confinement with no-slip boundary conditions and opening semi-angle $\alpha=\pi/6$. The rotlet is located at $\beta=\alpha/2$, see Fig. 1, at $z=0$ and points along the direction of the edge of the wedge (towards the reader). Results are presented in the radial–azimuthal plane, that is, in the plane normal to the edge of the wedge in (a) at height $z/\rho=0.1$ and (b) height $z/\rho=1$ above the plane of torque application. Conversely, the bottom row depicts the radial–axial plane containing the edge of the wedge for (c) $\theta=0$ and (d) $\theta=-\beta$. The depicted scaled velocity field is defined as $\bm{v}^{*}=\left(\rho^{2}/q_{\parallel}\right)\bm{v}$.
@@ III.2.2 Rotlet oriented in the transverse direction
Figure 3: Streamlines and quiver plots of the viscous flow field in analogy to Fig. 2, yet now induced by a radially oriented rotlet singularity pointing away from the tip of the wedge-like confinement. All other parameters are identical to those in Fig. 2.
Figure 4: Similarly to Fig. 2, streamlines and quiver plots of the viscous flow field are evaluated, yet now induced by an azimuthally oriented rotlet singularity, pointing in a direction perpendicular to the edge of the wedge and to the radial direction. All other parameters are identical to those in Fig. 2.
@@
In addition, $H_{z}=\Delta_{w}$ and $H_{z}^{\dagger}=\Delta_{w}^{\dagger}$.
@@
where, again, $\sigma=-1$ for $j=x$ and $\sigma=1$ for $j=y$, and the expression of $\Gamma_{\pm}^{-1}$ is given by Eq. (31). Defining $C=\cos(\beta+\delta)$ and $S=\sin(\beta+\delta)$, we obtain
@@ III.3 Solution in real space
$\psi_{j}(\theta,\nu)=q_{l}\left(\Lambda_{j}\operatorname{sh}(\theta\nu)+\Lambda_{j}^{\dagger}\operatorname{ch}(\theta\nu)\right),\qquad\xi_{j}(\theta,\nu)=q_{l}\left(H_{j}\operatorname{sh}(\theta\nu)+H_{j}^{\dagger}\operatorname{ch}(\theta\nu)\right).$ (47)
@@
see Gradshteyn and Ryzhik [43, p. 719], Eq. 6.672, or Prudnikov et al. [44, p. 390], Eq. 2.16.36(2). Here, $P_{n}$ is the Legendre function of the first kind of degree $n$, with argument
@@
where the free-space velocity $v_{j}^{\infty}$ is defined in Eq. (2). For axial point torque, we have $\phi_{z}=0$ and $\xi_{w}=0$. The expressions for the operator $\bm{\Psi}$ are obtained by substituting the integral forms given in Eq. (45) into the Papkovich–Neuber–Imai representation in Eq. (6), yielding
@@
Based on these expressions, from Eq. (52), we can thus calculate the resulting flow field within the wedge-shaped confinement for any orientation of the applied rotlet. In Fig. 2, we show a representative flow field generated by a torque singularity oriented in the axial direction, parallel to the edge of the wedge. In this example, the wedge has a semi-opening angle of $\alpha=\pi/6$, and the rotlet is positioned at $\beta=\alpha/2$. Near the singularity, the flow resembles that of an unbounded fluid. At greater distances, however, this unbounded solution is distorted by the wedge-shaped confinement. The location of the velocity maximum in planes that do not contain the torque singularity shifts away from the wedge surface as the height above the rotlet increases. In addition to the primary swirl induced by the rotlet, secondary swirling structures are present, though their velocities are significantly smaller.
@@ IV Solution in the planar-wall limit
The problem of a point torque singularity acting near a planar no-slip wall can be recovered in the limit of a wedge semi-opening angle $\alpha=\pi/2$. Results for this geometry have previously been reported, for example, by Spagnolie and Lauga [17]. In that case, the coefficients take particularly simple forms. They are summarized in Tab. 2 for both axial and transverse torques.
@@
$H_{x}$ $3\sin\beta\operatorname{sh}(\beta\nu)$
$H_{x}^{\dagger}$ $3\sin\beta\operatorname{ch}(\beta\nu)$
$H_{y}$ $\cos\beta\operatorname{sh}(\beta\nu)$
$H_{y}^{\dagger}$ $\cos\beta\operatorname{ch}(\beta\nu)$
$H_{z}$ $\operatorname{sh}(\beta\nu)\sin\delta$
$H_{z}^{\dagger}$ $\operatorname{ch}(\beta\nu)\sin\delta$
@@ IV.1 Axial torque
Using the integral presented in Appendix C of Ref. 15,
@@
where $\overline{s}=\mathscr{L}\left(b=\theta+\beta\right)=\sqrt{r^{2}+\rho^{2}+2\rho r\cos(\theta+\beta)+z^{2}}$, representing the value of $s$ with $\theta$ substituted by $\pi-\theta$, the image position relative to the wall. The resulting expressions for the velocity field fully agree with those reported by Spagnolie and Lauga [17].

### v2

@@ Abstract
Wedge-shaped geometries in low-Reynolds-number flows are of increasing importance, for instance, in the design of microfluidic devices. The corresponding Green’s functions describing the induced flow in response to a locally applied force were derived some time ago. To achieve a complete characterization of particle motion at low Reynolds numbers, we derive the flow response to locally applied torques. This is accomplished through a direct calculation based on the Fourier–Kontorovich–Lebedev transform using the Papkovich–Neuber representation of the hydrodynamic fields. We then illustrate the resulting flow fields, highlighting their structure, key features, and dependence on the geometry and orientation of the applied torque. Based on these solutions, we compute the corresponding hydrodynamic mobility tensor that couples torque and motion. Owing to the broken spatial symmetry imposed by the wedge-shaped confinement, a particle subjected to a torque will experience not only rotational motion but also translational motion. These results provide analytical tools relevant for predicting and controlling particle behavior in confined microfluidic environments.
@@ I Introduction
Wedge-shaped geometries have been studied in hydrodynamics because of their peculiar effect on intrinsic fluid flows, and because of their practical abundance. If a net velocity is prescribed on one surface of the wedge in radial direction under low Reynolds number conditions, vortices will be induced near the edge of the wedge, specifically at sufficiently low opening angles [1]. For instance, such net motion along one surface can be introduced by radial motion of the surface under no-slip conditions. Wedge-shaped geometries were further investigated under low-Reynolds-number conditions concerning their induced flows due to chemically active surfaces. It was demonstrated that resulting chemical concentration gradients can induce corresponding vortices near the edge of the wedge [2]. More recently, the self-diffusiophoretic motion of a catalytically active spherical particle confined within a wedge-shaped domain has been examined [3].
@@
Generally, the response to applied higher-order localized force distributions, such as force dipoles or quadrupoles, can be derived from this Green’s function, which is particularly relevant for modeling microswimmers and other active particles [18, 19, 20, 21, 22]. Application of a torque at one point within the fluid is another mode of action on the system. Here, we derive the resulting fluid flow of such a “rotlet” applied to the fluid explicitly. In practice, example situations of imposed rotations at small length scales comprise induced mixing by rotational components [23, 24] or micro-rheological investigations by rotating local probes [25, 26, 27, 28, 29, 30]. Typically, such rotations are induced using magnetic probe particles in reorienting external magnetic fields [31, 32]. In our context of the Green’s function, we would consider them as point-like.
Point torques (rotlets) have broad applications in microhydrodynamics. They provide a simple model for flows generated by beating cilia near surfaces, enabling predictions of fluid transport and mixing in ciliary carpets [33, 34]. In active systems such as colloidal microrollers, rotlets capture torque-driven flows near walls. They allow the analysis of particle transport, enhanced mixing, and collective interactions [35, 36, 37]. Similarly, bacterial hydrodynamics can be modeled using rotlets to describe the flows induced by rotating flagella or cell bodies. This strategy sheds light on surface accumulation, circular trajectories, and microscale mixing in bacterial suspensions [38, 39, 40].
@@ II.1 Governing equations
A general solution to Eq. (1) is given in the Papkovich–Neuber representation as [42, *Papkovich1932french, 44, 45]
@@ II.2.1 General overview and definitions
To determine the velocity field generated by a point torque in a wedge-shaped geometry, we employ the Fourier–Kontorovich–Lebedev (FKL) transform. In this approach, the axial $z$, and radial $r$ coordinates, are transformed into their corresponding wavenumbers, denoted by $k$ and $\nu$, respectively. The transform was first introduced by the Soviet mathematicians Kontorovich and Lebedev to address some boundary value problems [46, 47, 48]. The mathematical framework was later developed by Lebedev [49, 50]. We refer the reader to Erdélyi et al. [51, p. 75] for additional details on this integral transform.
@@
Generally, the FKL transform can be written as a double infinite integral over $r$ and $z$. Typically, there is no prescribed order of integration. In practice, however, it is common to first perform the forward Fourier transform, followed by the forward Kontorovich–Lebedev transform.
@@ II.2.2 Key Properties of the FKL Transform
Besides, many previous studies employed Mellin transforms to solve boundary-value problems in wedge-shaped domains [53, 54]. The Mellin transform is particularly effective when describing flows in two-dimensional corners, as outlined by Moffatt [1, 55]. For example, recently, this transform has been used to analyze flows induced in a two-dimensional corner by the chemical activity of the confining boundaries [2]. Both the Mellin transform and the here-employed Kontorovich–Lebedev transform can be applied to the radial coordinate in the governing equations. However, the Mellin transform is often more suitable for two-dimensional problems, while the KL transform is often better suited for three-dimensional problems. These transforms reduce the Laplace equation to an ordinary differential equation in the polar angle. In the three-dimensional case, the Fourier–Kontorovich–Lebedev transform is obtained by additionally applying a Fourier transform.
@@ III.1 Free-space solution
$f$ $\mathscr{T}_{i\nu}\{f\}=\int_{-\infty}^{\infty}\mathrm{d}z\,e^{ikz}\int_{0}^{\infty}f(r,z)K_{i\nu}(|k|r)\,r^{-1}\,\mathrm{d}r$
@@
We next employ the Table of Integral Transforms by Erdélyi et al. [56, p. 175] to derive the FKL transform of $1/s$ as
@@ III.2 Complementary solution
$A_{j}=\frac{2\pi q_{l}}{\nu\operatorname{sh}(\pi\nu)}\left(\frac{\Lambda_{j}}{\rho}+\mathrm{H}_{j}\,\frac{\partial}{\partial\rho}\right)K_{i\nu}(|k|\rho)\,,\qquad A_{j}^{\dagger}=\frac{2\pi q_{l}}{\nu\operatorname{sh}(\pi\nu)}\left(\frac{\Lambda_{j}^{\dagger}}{\rho}+\mathrm{H}_{j}^{\dagger}\,\frac{\partial}{\partial\rho}\right)K_{i\nu}(|k|\rho)\,$, (24)
where $l\in\{\parallel,\perp\}$ denoting the axial or transverse components of the point torque given in Eq. (4). Here, the coefficients $\Lambda_{j}$, $\mathrm{H}_{j}$, $\Lambda_{j}^{\dagger}$, and $\mathrm{H}_{j}^{\dagger}$ are functions of $\nu$, $\alpha$, and $\beta$, and, in the case of transverse torque, also depend on $\delta$.
For transverse torque, and for $j\in\{x,y,w\}$, the coefficients are expressed as
@@
which is expressed in terms of $\Delta_{j}$ and $\Delta_{j}^{\dagger}$. They are functions of $\nu$, $\alpha$, $\beta$, and $\delta$. By expressing the coefficients in this form, the dependence of $A_{j}$ and $A_{j}^{\dagger}$ on $k$ becomes explicit. This step facilitates integration with respect to $k$ during the inverse transforms. The reason for including $\nu\operatorname{sh}(\pi)$ in the denominator of these expressions is to simplify the resulting formulas upon performing the inverse transformation; see Eq. (13).
@@ III.2.1 Rotlet oriented in the axial direction
In addition, $\mathrm{H}_{w}=\mathrm{H}_{w}^{\dagger}=0$. The expressions for the remaining coefficients are considerably more complex, so we simplify them using abbreviations. We define the shorthand notations
@@
$\Lambda_{j}$ $=\Gamma_{\sigma}\,\nu\left(a_{j}\nu\sin(2\alpha)+b_{j}\right)\,,\qquad\quad\mathrm{H}_{j}=\Gamma_{\sigma}\left(p_{j}\nu\sin(2\alpha)+w_{j}\right)\,$, (32a)
$\Lambda_{j}^{\dagger}$ $=\Gamma_{-\sigma}\,\nu\left(a_{j}^{\dagger}\nu\sin(2\alpha)+b_{j}^{\dagger}\right)\,,\quad\quad\mathrm{H}_{j}^{\dagger}=\Gamma_{-\sigma}\left(p_{j}^{\dagger}\nu\sin(2\alpha)+w_{j}^{\dagger}\right)\,$, (32b)
where $\sigma=$- for $j=x$ and $\sigma=+$ for $j=y$. The coefficients defining $\Lambda_{j}$ and $\Lambda_{j}^{\dagger}$ are obtained as
@@
The parameters defining $\mathrm{H}_{j}$ and $\mathrm{H}_{j}^{\dagger}$ are obtained as
@@
Figure 2: Streamlines and quiver plots of the viscous flow field induced by an axially oriented rotlet singularity in a wedge-like confinement with no-slip boundary conditions and semi–opening angle $\alpha=\pi/6$. The rotlet is located at $\beta=\alpha/2$, see Fig. 1, at $z=0$ and points along the direction of the edge of the wedge (towards the reader). Results are presented in the radial–azimuthal plane, that is, in the plane normal to the edge of the wedge in (a) at height $z/\rho=0.1$ and (b) height $z/\rho=1$ above the plane of torque application. Conversely, the bottom row depicts the radial–axial plane containing the edge of the wedge for (c) $\theta=0$ and (d) $\theta=-\beta$. The depicted scaled velocity field is defined as $\bm{v}^{*}=\left(\rho^{2}/q_{\parallel}\right)\bm{v}$, while the plots are generated using the results derived in Sec. III.2.1.
@@ III.2.2 Rotlet oriented in the transverse direction
Figure 3: Streamlines and quiver plots of the viscous flow field in analogy to Fig. 2, yet now induced by a radially oriented rotlet singularity pointing away from the tip of the wedge-like confinement. All other parameters are identical to those in Fig. 2, while these plots are generated from the results in Sec. III.2.2 with $\delta=0$.
Figure 4: Similarly to Fig. 2, streamlines and quiver plots of the viscous flow field are evaluated, yet now induced by an azimuthally oriented rotlet singularity, pointing in a direction perpendicular to the edge of the wedge and to the radial direction. All other parameters are identical to those in Fig. 2, whereas the present plots are generated from the results in Sec. III.2.2 with $\delta=\pi/2$.
@@
In addition, $\mathrm{H}_{z}=\Delta_{w}$ and $\mathrm{H}_{z}^{\dagger}=\Delta_{w}^{\dagger}$.
@@
where, again, $\sigma=$- for $j=x$ and $\sigma=+$ for $j=y$, and the expression of $\Gamma_{\pm}^{-1}$ is given by Eq. (31). Defining $C=\cos(\beta+\delta)$ and $S=\sin(\beta+\delta)$, we obtain
@@ III.3 Solution in real space
$\psi_{j}(\theta,\nu)=q_{l}\left(\Lambda_{j}\operatorname{sh}(\theta\nu)+\Lambda_{j}^{\dagger}\operatorname{ch}(\theta\nu)\right),\qquad\xi_{j}(\theta,\nu)=q_{l}\left(\mathrm{H}_{j}\operatorname{sh}(\theta\nu)+\mathrm{H}_{j}^{\dagger}\operatorname{ch}(\theta\nu)\right).$ (47)
@@
see Gradshteyn and Ryzhik [57, p. 719], Eq. 6.672, or Prudnikov et al. [58, p. 390], Eq. 2.16.36(2). Here, $P_{n}$ is the Legendre function of the first kind of degree $n$, with argument
@@
where the free-space velocity $v_{j}^{\infty}$ is defined in Eq. (2). For axial point torque, we have $\phi_{z}=0$ and $\xi_{w}=0$. The expressions for the operator $\bm{\Psi}$ are obtained by substituting the integral forms given in Eq. (45) into the Papkovich–Neuber representation in Eq. (6), yielding
@@
The infinite integrals in Eq. (52) are evaluated numerically using the built-in command NIntegrate in the computer algebra system Mathematica [59]. Since the integrands involve several complicated functions, a sufficiently large working precision is required to ensure accurate results. In practice, a working precision of 100 is found to provide a reasonable balance between accuracy and computational efficiency. As the integrands decay rapidly with increasing $\nu$, the integrals are truncated at a finite upper limit, which we take to be 100. Larger values slow down the numerical integration without notably affecting the results, as far as we observed. Mathematica scripts used to generate these flow fields are freely available on Zenodo [60].
Based on these expressions, from Eq. (52), we can thus calculate the resulting flow field within the wedge-shaped confinement for any orientation of the applied rotlet. In Fig. 2, we show a representative flow field generated by a torque singularity oriented in the axial direction, parallel to the edge of the wedge. In this example, the wedge has a semi–opening angle of $\alpha=\pi/6$, and the rotlet is positioned at $\beta=\alpha/2$. Near the singularity, the flow resembles that of an unbounded fluid. At greater distances, however, this unbounded solution is distorted by the wedge-shaped confinement. The location of the velocity maximum in planes that do not contain the torque singularity shifts away from the wedge surface as the height above the rotlet increases. In addition to the primary swirl induced by the rotlet, secondary swirling structures are present, though their velocities are significantly smaller.
@@ IV Solution in the planar-wall limit
The problem of a point torque singularity acting near a planar no-slip wall can be recovered in the limit of a wedge semi–opening angle $\alpha=\pi/2$. Results for this geometry have previously been reported, for example, by Blake and Chwang [61]; see also Spagnolie and Lauga [18]. In that case, the coefficients take particularly simple forms. They are summarized in Tab. 2 for both axial and transverse torques.
@@
$\mathrm{H}_{x}$ $3\sin\beta\operatorname{sh}(\beta\nu)$
$\mathrm{H}_{x}^{\dagger}$ $3\sin\beta\operatorname{ch}(\beta\nu)$
$\mathrm{H}_{y}$ $\cos\beta\operatorname{sh}(\beta\nu)$
$\mathrm{H}_{y}^{\dagger}$ $\cos\beta\operatorname{ch}(\beta\nu)$
$\mathrm{H}_{z}$ $\operatorname{sh}(\beta\nu)\sin\delta$
$\mathrm{H}_{z}^{\dagger}$ $\operatorname{ch}(\beta\nu)\sin\delta$
@@ IV.1 Axial torque
Using the integral presented in Appendix C of Ref. 16,
@@
where $\overline{s}=\mathscr{L}\left(b=\theta+\beta\right)=\sqrt{r^{2}+\rho^{2}+2\rho r\cos(\theta+\beta)+z^{2}}$, representing the value of $s$ with $\beta$ substituted by $\pi-\beta$, the image position relative to the wall.
Using the expressions for the harmonic functions given in Eqs. (60), the components of the velocity field are first obtained in cylindrical coordinates. Upon transforming to Cartesian coordinates, the resulting velocity field is found to be in full agreement with the results reported by Blake and Chwang [61].